OpenStudy (anonymous):

What is the value of f[g(5)] for the functions f(x) = 2x + 1 and g(x) = 2x - 5 ? Please I don't get it at all.

OpenStudy (anonymous):

whats g(5)?

OpenStudy (anonymous):

so first solve function g(x) with x = 5 then whatever you get from that, substitute into f(x) and solve that

OpenStudy (anonymous):

I have no idea, im taking a test and thats the problem they gave me

OpenStudy (anonymous):

uh...

OpenStudy (anonymous):

g(5)=2(5)-5=10-5=5

OpenStudy (anonymous):

so five is the answer?

OpenStudy (anonymous):

no 5 is g(5):) now f(g(5))=f(5)=2(5)+1=11 which is the answer

OpenStudy (anonymous):

OHHH i see how you did it, thanks!

OpenStudy (anonymous):

good

OpenStudy (amistre64):

"im taking a test" is always rife with bad vibes ...